# An Evolved Universal Transformer Memory

### v2

@@ 1 Introduction
We share our full training code to facilitate future advances in foundation models through memory.
@@ 3.1 Attention spectrograms for model-agnostic feature extraction
The feature extraction framework of NAMMs is designed to be agnostic to the parameterization of the base transformer they are applied for. In particular, we build a representation for each token in the current KV cache memory from its corresponding column vector in the attention matrix . To meaningfully compress this unbounded vector signal, we process it via an STFT with a fixed-sized Hann window (Figure 2, left). This operation produces a spectrogram representation of the attention columns , representing the frequencies with how the queries attend to each of the stored key tokens (indexed by ) on a compressed time-axis (indexed by ). Thus, this representation exposes precisely the knowledge of how each token’s relative importance varies across all past queries in a compact form factor, discarding all other information specific to the learned transformer weights.
@@ 3.2 Memory model design and cross-token communication
3.3 Training NAMMs with incremental evolution
We evolve the network weights of our NAMMs to directly optimize the performance on a subset of long-context language modeling tasks from LongBench (Bai et al., 2023). As we share a single across all layers, even with our largest NAMM we only evolve about 4000 total parameters. We use the seminal CMA-ES optimization algorithm (Hansen, 2006) and apply NAMM atop a Llama 3 8B base model (Dubey et al., 2024) with a context extended from 8192 to 32768 tokens via NTK-aware positional interpolation (bloc97, 2023). Due to the inference costs of LMs with long inputs, we sample a subset of different prompts from each task in each generation and propose training in an incremental fashion: starting from a single task, and adding additional tasks at later training stages. Empirically, we found both these choices to provide effective regularization, improving generalization (see Appendix C). The performance of modern LMs on LongBench varies considerably across tasks, and even across different task prompts. Hence, instead of using the raw scores, we opt to maximize normalized performance relative to the vanilla base model’s stored evaluation performance on each same subset of prompts, retaining all tokens in its KV cache memory.
We choose three tasks from different LongBench categories across both English and Chinese where the Llama 3 base model seems to particularly struggle: PassageRetrieval-en, DuReader, and NarrativeQA. We evolve our NAMM for 300 generations in its first incremental phase, 250 in its second, and 120 in its third. We diminish the number of generations to counteract the increasing costs with each additional phase and make more efficient use of computational resources. At the end of each phase, we resume from the best previous checkpoint. We provide training curves of our main backward-attention model in Figure 4, showing the average and standard deviation of the normalized batch performance across the population (left), together with the normalized per-task and average performance on all samples of the optimized mean from CMA-ES (right). We refer to Appendix A for additional architectural and optimization details, together with the set of hyper-parameters. We also provide additional statistics and training curves for other memory model designs in Appendix C.
@@ 4 Experimental Results
In this section, we evaluate and analyze evolved NAMMs as compared to full-context transformers and two recent hand-designed methods for KV cache management: H2O (Zhang et al., 2024c) and L2 (Devoto et al., 2024). We compare each method in terms of absolute and normalized performance and also provide the resulting average cache size recorded at the end of each prompt. We first consider three long-context language modeling benchmarks spanning 36 diverse tasks in three languages, using the same Llama 3 8B base transformer from training. Then, we evaluate the capabilities of zero-shot transferring NAMMs to other unseen transformers and task domains. In particular, we not only consider transfer to larger LMs, but also transformers with tokens constructed from modalities other than language. Across all these settings, we also compare BAM with a simpler 2-layer MLP architecture and provide summarized results after every stage of incremental evolutions. We refer to Appendix C additional evaluations, ablation studies, and all learning curves. Lastly, we perform targeted qualitative analysis to understand the behavior of our new memory framework.
@@ 4.1 Long-context language understanding
| H2O | 73.00 (1.00) | 90.03 (1.01) | 46.48 (1.00) | 34.00 (0.85) | 2.18 (1.47) | 9.93 (0.82) | 27.76 (0.96) | 69.37 (1.00) | 65.44 (1.00) | 28.37 (0.99) | N/A | 8192 (0.25) |
| L2 | 66.41 (0.91) | 84.92 (0.95) | 45.78 (0.98) | 34.38 (0.86) | 3.13 (2.11) | 11.00 (0.90) | 28.68 (1.00) | 73.45 (1.06) | 55.20 (0.85) | 27.42 (1.00) | N/A | 8192 (0.25) |
| NAMM (Ours) | 73.00 (1.00) | 89.81 (1.00) | 46.35 (1.00) | 40.00 (1.00) | 3.04 (2.05) | 27.55 (2.26) | 28.60 (0.99) | 69.53 (1.01) | 66.35 (1.02) | 29.33 (1.11) | 1.07 | 8155 (0.25) |
Longbench. In Table 2, we provide results across all LongBench tasks (Bai et al., 2023). Our NAMM yields concrete improvements to the Llama 3 8B transformer both when considering the full set or exclusively the held-out set of test tasks that were not used for evolution, with improvements of 11% and 7% respectively. At the same time, our NAMM also yields efficiency side benefits, notably reducing the context-extended KV cache size. Instead, while both H2O and L2 produce even smaller cache sizes, they both come with some performance costs - in line with their stated objective of retaining rather than improving the original full-context performance. These results indicate how two different hand-designed extremes, either retaining all tokens or aggressively dropping them, come with their own set of downsides for either efficiency or performance. On the other hand, NAMMs are the only methods making consistent improvements from the base model across both axes, highlighting how end-to-end evolutionary optimization can open new orthogonal directions beyond what is feasible with manually-designed heuristics.
@@
ChouBun. Our new benchmark focuses on tasks designed exclusively in Japanese, a novel language unseen during NAMMs training. We hope this benchmark might itself be a valuable contribution to the research community, allowing the assessment of long-context capabilities in multilingual LLMs beyond the already-ubiquitous English and Chinese. We provide further benchmark statistics, details about task composition, together with evaluation metrics for a wider range of popular LLMs in Appendix B. In Table 4, we report our results evaluating NAMMs. Once again, we observe a clear contrast with prior hand-designed methods. While integrating either H2O or L2 leads to notable performance drops, NAMMs provides substantial improvements, with overall performance up by 15% from the full-context Llama 3 8B base model.
@@ 4.2 Zero-shot transfer across architectures and modalities
Reinforcement learning. In Table 7, we provide our zero-shot transfer results for the offline reinforcement learning setting, where we apply NAMMs atop a decision transformer (Chen et al., 2021b) and consider the canonical continuous-control tasks from the D4RL benchmark (Fu et al., 2020). We find our NAMM improves the base transformer quite considerably in this domain across eight out of nine offline tasks with over 9% overall gains, opposing the performance loss of the other efficient baselines. We posit that since the nature of the decision transformer optimization is closely tied to behavior cloning, the ability to discard part of the context is likely to allow NAMMs to forget and avoid imitating previous mistakes autoregressively. In support of this hypothesis, we observed slightly higher average rewards in the transitions for the retained tokens (by 1.4%, 0.8%, and 12.3% for the Hopper, Walker2d, and HalfCheetah environments, respectively).
@@ 4.3 Understanding Neural Attention Memory Models
Backward attention cross-token interactions. To analyze cross-token interactions learned through our BAM architecture, we record the gradients of each token score with respect to all input features for all tokens in memory after storing 1024 tokens, i.e., for . We denote these quantities as . We provide a qualitative visualization of our results on the PassageRetrieval-en task for a randomly selected layer and prompt in Figure 7. On the left subplot, we provide a visualization of the squared magnitudes for each combination of tokens (either scored or attended upon in BAM, i.e., indexed by or ). Here, the effects of the backward mask are clearly visible, allowing tokens to exclusively attend to later ones, where . Predictably, these magnitudes mostly peak on the subplot’s diagonal, indicating the self-influence that each token’s features have on its corresponding output score. However, there are also notable exceptions, as shown in the center subplot, where we overlap three slices from our left surface plot corresponding to the gradients of the first, together with the highest and lowest-scored tokens in memory (respectively indexed by 0, 292, and 800). We provide additional directional information of each gradient vector from these slices in the right subplot, where we take its dot product with the scored token’s own feature vector . After the first notable spike, at , most other dot-product spikes with the largest magnitudes consistently have negative values. Hence we can logically deduce that the scores of these tokens would benefit from pushing the representations of future tokens away from their own. This result appears to validate the hypothesis that BAM learns a mechanism for cross-token competition, incentivizing diversity and promoting tokens covering unique frequencies in the attention spectrogram.
Additional analysis. We provide additional analytic results in Appendix D. For instance, we compare the generated responses before and after the introduction of NAMMs in a very long context task, and show the sensitivities of the token scores for each input feature. These results illustrate how NAMMs learn to overcome different failure modes of long context LMs, and that they mainly rely upon high-frequency components of the attention matrices, further evidencing the need to go beyond simple strategies and the potential of end-to-end learning for token-level memory systems.
@@ 5 Related works
Devoto et al. (2024) and Yao et al. (2024) try to identify the least important tokens to evict using heuristics such as L2 magnitude and entropy. Alternative strategies include considering simple statistics from the attention matrix (Liu et al., 2024; Oren et al., 2024; Zhang et al., 2024c). Ge et al. (2024) and Li et al. (2024b) build on these ideas by applying multiple strategies based on matching specific attention patterns. Unlike this prior work, our approach uniquely employs a black-box model to learn token-level memory management and shows potential for improving both the performance and efficiency of transformers. We refer to App. E for references and connections to the wider literature, including efficient architectures, memory, and evolution.
@@ 6 Discussion and future work
This work introduced Neural Attention Memory Models, providing a new framework to enhance the performance of transformers while significantly reducing memory footprint. By evolving NAMMs on top of pre-trained LMs, we demonstrated their effectiveness across diverse long-context tasks in three languages, significantly surpassing previous hand-designed KV cache eviction frequently hindering performance, and the original model relying on costly full-context conditioning. Our carefully designed approach also enabled NAMMs, trained solely on language tasks, to achieve zero-shot transferability across architectures, input modalities, and task domains. This work has only begun to explore the design space of our memory models, which we anticipate might offer many new opportunities to advance future generations of transformers. In this regard, we believe NAMMs should not be viewed as a replacement for gradient-based optimization, but rather an orthogonal framework that could be combined and alternated with parameter fine-tuning. Such an extension has the potential to unlock efficient long-context training, drawing parallels to the iterative process of learning and evolution that shaped human memory.
@@ A.2 Feature extraction and architecture details
where we use to denote the EMA’s discount factor. To expedite learning the weights of our architecture, we ensure all spectrogram features have unit variance at initialization across our training data, using the statistics of the base Llama 3 model computed on the first task employed in incremental learning (PassageRetrieval). Finally, we also concatenate a small eight-dimensional sinusoidal positional embedding using the oldness of each token, i.e., the amounts of new queries observed since its introduction in the KV cache.
@@ A.3 Zero-shot transfer
For our zero-shot transfer experiments, we consider a Llama 3 transformer with 70B parameters (Dubey et al., 2024), a Llava Next Video transformer with 7B parameters (Zhang et al., 2024b), and a decision transformer (Chen et al., 2021b) with about 1M parameters. For our 70B experiments, we follow the exact same setup as when evaluating our 7B Llama model used in training. For our video-language model, we extract image tokens from 48 uniformly sampled frames, 6912 in total. We also slightly shift the selection score threshold by 5, to counteract the lower number of total tokens and get a comparable average cache size to the L2 and H2O baselines. We adapt the code and follow the standardized experimental setup from Li et al. (2024a). For the reinforcement learning experiments, we encode each state, action, and return-to-go into separate tokens and do not apply any restrictions or modifications to our standard NAMM LM setup. We average the performance collected over 20 random seeds to account for the stochasticity of the initial state in the Gym Mujoco environments (Brockman et al., 2016). Our RL experiments adapt the checkpoints and setup provided by Beeching & Simonini (2022).
Appendix B ChouBun details

### v3

@@ 3.1 Attention spectrograms for model-agnostic feature extraction
The feature extraction framework of NAMMs is designed to be agnostic to the parameterization of the base transformer they are applied for. In particular, we build a representation for each token in the current KV cache memory directly from its corresponding unmodified column vector in the attention matrix . To meaningfully compress this unbounded vector signal, we process it via an STFT with a fixed-sized Hann window (Figure 2, left). This operation produces a spectrogram representation of the attention columns , representing the frequencies with how the queries attend to each of the stored key tokens (indexed by ) on a compressed time-axis (indexed by ). Thus, this representation exposes precisely the knowledge of how each token’s relative importance varies across all past queries in a compact form factor, discarding all other information specific to the learned transformer weights.
@@ 3.2 Memory model design and cross-token communication
In practice, when applying NAMMs, we only affect the KV cache of the base model with a fixed frequency, once every steps. When feeding longer prompts to our model, we simply split the tokens into -sized chunks. We summarize the full execution pipeline of NAMMs in Algorithm 1. We refer to Appendix A and our shared code for additional implementation details and discussion.
3.3 Incremental evolution
We evolve the network weights of our NAMMs to directly optimize the performance on a subset of long-context language modeling tasks from LongBench (Bai et al., 2023). As we share a single across all layers, even with our largest NAMM we only evolve about 4000 total parameters. We use the seminal CMA-ES optimization algorithm (Hansen, 2006) and apply NAMM atop a Llama 3 8B base model (Dubey et al., 2024) with a context extended from 8192 to 32768 tokens via NTK-aware positional interpolation (bloc97, 2023). Due to the inference costs of LMs with long inputs, we sample a subset of different prompts from each task in each generation and propose training in an incremental fashion: starting from a single task, and adding additional tasks at later training stages. Empirically, we found both these choices to provide effective regularization, improving generalization (see Appendix C). The performance of modern LMs on LongBench varies considerably across tasks, and even across different task prompts. Hence, instead of using the raw scores, we opt to maximize normalized performance relative to the vanilla base model’s stored evaluation performance on each same subset of prompts, retaining all tokens in its KV cache memory. Using evolution, we note that our training loop simply corresponds to running inference NAMMs atop the base, requiring no expensive backpropagation or dedicated hardware.
We choose three tasks from different LongBench categories across both English and Chinese where the Llama 3 base model seems to particularly struggle: PassageRetrieval-en, DuReader, and NarrativeQA; optimizing the normalized exact match, ROUGE-L, and F1 metrics, respectively. We evolve our NAMM for 300 generations in its first incremental phase, 250 in its second, and 120 in its third. We diminish the number of generations to counteract the increasing costs with each additional phase and make more efficient use of computational resources. At the end of each phase, we resume from the best previous checkpoint. We provide training curves of our main backward-attention model in Figure 4, showing the average and standard deviation of the normalized batch performance across the population (left), together with the normalized per-task and average performance on all samples of the optimized mean from CMA-ES (right). We refer to Appendix A for additional architectural and optimization details, together with the set of hyper-parameters. We also provide additional statistics and training curves for other memory model designs in Appendix C.
@@ 4 Experimental Results
In this section, we evaluate and analyze evolved NAMMs as compared to full-context transformers and three recent hand-designed methods for KV cache management: H2O (Zhang et al., 2024c) and L2 (Devoto et al., 2024), and FastGen (Ge et al., 2024). We compare each method in terms of absolute and normalized performance and also provide the resulting average cache size recorded at the end of each prompt. We first consider three long-context language modeling benchmarks spanning 36 diverse tasks in three languages, using the same Llama 3 8B base transformer from training. Then, we evaluate the capabilities of zero-shot transferring NAMMs to other unseen transformers and task domains. In particular, we not only consider transfer to larger LMs, but also transformers with tokens constructed from modalities other than language. Across all these settings, we also compare BAM with a simpler 2-layer MLP architecture and provide summarized results after every stage of incremental evolutions. We refer to Appendix C additional evaluations (e.g., transferring NAMMs to a Mistral LM), ablation studies (e.g., comparing different architectures and input features), and all learning curves. Lastly, we conclude the Section with a targeted qualitative analysis, aimed at understanding the behavior of our new memory framework.
@@ 4.1 Long-context language understanding
| Base model | 73.00 (1.00) | 89.45 (1.00) | 46.54 (1.00) | 40.00 (1.00) | 1.48 (1.00) | 12.18 (1.00) | 28.80 (1.00) | 69.09 (1.00) | 65.17 (1.00) | 28.86 (1.00) | N/A | 10107 (1.00) |
| H2O | 73.00 (1.00) | 90.03 (1.01) | 46.48 (1.00) | 34.00 (0.85) | 2.18 (1.47) | 9.93 (0.82) | 27.76 (0.96) | 69.37 (1.00) | 65.44 (1.00) | 28.37 (0.99) | N/A | 6662 (0.66) |
| L2 | 66.41 (0.91) | 84.92 (0.95) | 45.78 (0.98) | 34.38 (0.86) | 3.13 (2.11) | 11.00 (0.90) | 28.68 (1.00) | 73.45 (1.06) | 55.20 (0.85) | 27.42 (1.00) | N/A | 6662 (0.66) |
| FastGen | 73.00 (1.00) | 88.76 (0.99) | 46.40 (1.00) | 36.00 (0.90) | 1.15 (0.78) | 10.23 (0.84) | 27.01 (0.94) | 69.34 (1.00) | 64.50 (0.99) | 27.88 (0.95) | N/A | 9538 (0.94) |
| NAMM (Ours) | 73.00 (1.00) | 89.81 (1.00) | 46.35 (1.00) | 40.00 (1.00) | 3.04 (2.05) | 27.55 (2.26) | 28.60 (0.99) | 69.53 (1.01) | 66.35 (1.02) | 29.33 (1.11) | 1.07 | 8409 (0.83) |
Longbench. In Table 2, we provide results across all LongBench tasks (Bai et al., 2023) and in Figure 5 we provide a summarized comparison varying the maximum cache size of H2O and L2 (we provide a similar analysis for FastGen in Figure 9). Our NAMM yields concrete improvements to the Llama 3 8B transformer both when considering the full set or exclusively the held-out set of test tasks that were not used for evolution, with improvements of 11% and 7% respectively. At the same time, our NAMM also yields efficiency side benefits, notably reducing the context-extended KV cache size. Instead, H2O, L2, and Fastgen all come with performance costs which notably grow the smaller their cache sizes - in line with their stated objective of retaining rather than improving the original full-context performance. These results emphasize the inevitable tradeoff induced by prior hand-designed methods, able to obtain efficiency gains but at increasing performance costs due to their lossy heuristics. On the other hand, we find NAMMs successfully provide a paradigm shift, yielding consistent improvements from the base model across both performance and efficiency axes by learning to discard unhelpful information, highlighting how end-to-end evolutionary optimization can open new orthogonal directions beyond what is feasible with manually-designed heuristics.
@@
ChouBun. Our new benchmark focuses on tasks designed exclusively in Japanese, a novel language unseen during NAMMs training. We hope this benchmark might itself be a valuable contribution to the research community, allowing the assessment of long-context capabilities in multilingual LLMs beyond the already-ubiquitous English and Chinese. We provide further benchmark statistics, details about task composition, together with evaluation metrics for a wider range of popular LLMs in Appendix 10. In Table 4, we report our results evaluating NAMMs. Once again, we observe a clear contrast with prior hand-designed methods. While integrating either H2O or L2 leads to notable performance drops, NAMMs provides substantial improvements, with overall performance up by 15% from the full-context Llama 3 8B base model.
@@ 4.2 Zero-shot transfer across architectures and modalities
Reinforcement learning. In Table 7, we provide our zero-shot transfer results for the offline reinforcement learning setting, where we apply NAMMs atop a decision transformer (Chen et al., 2021b) using the open-sourced models from Beeching & Simonini (2022) pre-trained on the canonical the continuous-control tasks from D4RL (Fu et al., 2020). We find our NAMM improves the base transformer quite considerably in this domain across eight out of nine offline tasks with over 9% overall gains, opposing the performance loss of the other efficient baselines. We posit that since the nature of the decision transformer optimization is closely tied to behavior cloning, the ability to discard part of the context is likely to allow NAMMs to forget and avoid imitating previous mistakes autoregressively. In support of this hypothesis, we observed slightly higher average rewards in the transitions for the retained tokens (by 1.4%, 0.8%, and 12.3% for the Hopper, Walker2d, and HalfCheetah environments, respectively).
@@ 4.3 Understanding Neural Attention Memory Models
Additional analysis. We provide additional analytic results in Appendix D. For instance, we analyze how the presence of each token in memory affects the scores of the other tokens, we compare the generated responses before and after the introduction of NAMMs in a very long context task, and show the sensitivities of the token scores for each input feature. These results show that NAMMs learn mechanisms for ‘cross-token’ competition relying on high-frequency components of the attention matrices and illustrate how they learn to overcome different failure modes of long context LMs, further evidencing the need to go beyond simple strategies and the potential of end-to-end learning for token-level memory systems.
@@ 5 Related works
Devoto et al. (2024) and Yao et al. (2024) try to identify the least important tokens to evict using heuristics such as L2 magnitude and entropy to improve efficiency. Alternative strategies include considering simple statistics from the attention matrix (Liu et al., 2024b; Oren et al., 2024; Zhang et al., 2024c). Ge et al. (2024) and Li et al. (2024b) build on these ideas by applying multiple strategies based on matching specific attention patterns. Motivated by similar considerations, Nawrot et al. (2024) proposed directly fine-tuning the original base transformer to compress the KV cache while minimizing a regularization loss to preserve the original base model’s behavior and limit performance degradation. Complementary to our method, MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) propose merging attention heads during training to improve inference throughput. Similarly, also KV cache quantization is another orthogonal area where different hand-designed strategies have been proposed (Hooper et al., 2024; Dong et al., 2024a; b), with even recent work empirically showing their direct compatibility with token eviction methods (Liu et al., 2024a). We note that, unlike this prior work, our approach uniquely learns a black-box model to maximize performance through token-level memory management and shows potential for providing improvements to both the effectiveness and efficiency of transformers. We refer to App. E for references and to the wider literature, including efficient architectures, memory, and evolution.
@@ 6 Discussion and future work
This work introduced Neural Attention Memory Models, providing a new framework to enhance the performance of transformers while significantly reducing memory footprint. By evolving NAMMs on top of pre-trained LMs, we demonstrated their effectiveness across diverse long-context tasks in three languages, significantly surpassing previous hand-designed KV cache eviction frequently hindering performance, and the original model relying on costly full-context conditioning. Our carefully designed approach also enabled NAMMs, trained solely on language tasks, to achieve zero-shot transferability across architectures, input modalities, and task domains. While NAMMs do appear to provide benefits beyond what achieved with hand-designed strategies, we believe there is much room for improvement (e.g., see Limitations F). This work has only begun to explore the design space of our memory models, which we anticipate might offer many new opportunities to advance future generations of transformers. In this regard, we believe NAMMs should not be viewed as a replacement for gradient-based optimization, but rather an orthogonal framework that could be combined and alternated with parameter fine-tuning. Such an extension has the potential to unlock efficient long-context training, drawing parallels to the iterative process of learning and evolution that shaped human memory.
@@ A.2 Feature extraction and architecture details
where we use to denote the EMA’s discount factor. To expedite learning the weights of our architecture, we ensure all spectrogram features have unit variance at initialization across our training data, using the statistics of the base Llama 3 model computed on the first task employed in incremental learning (PassageRetrieval). Finally, we also concatenate a small eight-dimensional sinusoidal positional embedding using the oldness of each token, i.e., the amounts of new queries observed since its introduction in the KV cache. We provide an extended summarized pseudocode description of the execution pipeline in Algortihm 2 (to complement Algorithm 1 in the main text).
@@ A.3 Zero-shot transfer
For our zero-shot transfer experiments, we consider a Llama 3 transformer with 70B parameters (Dubey et al., 2024), a Llava Next Video transformer with 7B parameters (Zhang et al., 2024b), and a decision transformer (Chen et al., 2021b) with about 1M parameters. For our 70B experiments, we follow the exact same setup as when evaluating our 7B Llama model used in training. For our video-language model, we extract image tokens from 48 uniformly sampled frames, 6912 in total. We also slightly shift the selection score threshold by 5, to counteract the lower number of total tokens and get a comparable average cache size to the L2 and H2O baselines. We adapt the code and follow the standardized experimental setup from Li et al. (2024a). For the reinforcement learning experiments, we encode each state, action, and return-to-go into separate tokens and do not apply any restrictions or modifications to our standard NAMM LM setup. We average the performance collected over 20 random seeds to account for the stochasticity of the initial state in the Gym Mujoco environments (Brockman et al., 2016). Rather than re-training a decision transformer from scratch, our RL experiments adapt the open-sourced checkpoints and implementation provided by Beeching & Simonini (2022). We would like that note that on some task-dataset combinations of D4RL, these checkpoints appear to yield lower performance than what was reported in the original decision transformer paper (e.g., Walker pre-trained on medium-expert data) Chen et al. (2021b). However, we do not believe these differences should affect our conclusions as we used the same base model for all our memory management baselines.
A.4 Evolutionary optimization
As described in Section 3, we optimize NAMMs with the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen, 2006). Being an evolutionary algorithm, CMA-ES does not require any gradient information and can directly optimize black-box undifferentiable metrics. This property allows us to both optimize for the non-differentiable token selection task of our NAMMs and also maximize non-differentiable task performance metrics directly. In the case of the LongBench (Bai et al., 2023) tasks considered for training, these metrics correspond to exact match accuracy (PassageRetrieval-en), ROUGE-L score (DuReader), and F1 score (NarrativeQA).
On a high level, given a neural network with parameters, CMA-ES maintains a mean vector and a covariance matrix . Then, it repeats the following steps:
1.
Sampling. CMA-ES generates a population of neural networks, sampling their parameters from the multivariate normal distribution.
2.
Evaluation. Each population candidate is evaluated to the objective function for the objective function used.
3.
Updating. By both selecting a subset of the population candidates and also weighting them based on their overall ranking the mean and covariance are updated towards higher-performing regions of the search space.
We provide the main hyper-parameters in Table 9 and refer to either the work by Hansen (2006) or our shared code for the full implementation details.
A.5 FastGen implementation and tuning
We re-implemented the recent FastGen method proposed by Ge et al. (2024), which proposes to adopt a hand-designed combination of different strategies targeted to retain tokens with high attention values, belonging to recent words, or encoded from particular grammatical features (i.e., punctuation, ‘special tokens’). In particular, after observing the input prompt, FastGen performs a ‘profiling step’ where the strategy able to evict the most amount of tokens is selected such that:
Here, is the full-cache attention matrix, is the ‘reconstructed’ attention matrix re-calculated after performing a softmax between each layer’s queries and keys with masked-out entries for the keys evicted by the individual strategies. Furthermore, T is the main threshold hyper-parameter, determining how aggressively FastGen is allowed to prune tokens even if resulting in degradation to the attention-reconstruction heuristic.
We note that, unlike our other baselines, FastGen is only directly compatible with language modeling tasks. This is because one of the main ways it differs from H2O is by preserving particular grammar-based tokens in some of its strategies (e.g., punctuation, special words, etc.). Thus, as this baseline was specifically designed for LMs rather than arbitrary transformers, we did not consider applying it in the 0-shot transfer settings, and only focused on Llama 3 8B.
We note that as we are dealing with much longer prompts (sometimes far beyond tens/hundreds of thousand tokens), for efficiency consideration, we performed the profiling steps in our re-implementation after the first 4096 tokens any prompt exceeds this length. We also found to avoid losing too much performance over the base model on longer context tasks we had to retune its main ‘threshold.’ We selected T=0.999, as this choice allowed FastGen to retain over 95% normalized performance while still discarding a non-trivial portion of tokens on all LongBench, as shown in Figure 9. Other than the main threshold for attention reconstruction, FastGen has two other main hyperparameters: the ‘recency ratio,’ the ‘attention ratio’ determining the portion of most recent tokens or with the highest attention values to retain in its individual strategies. We set these hyper-parameters to 0.3, following the paper’s recommendation.
Appendix B Benchmark descriptions
B.1 ChouBun details